# Dissipative spatiotemporal soliton in a driven waveguide laser

### v1

@@ 3 Discussion and Conclusion
Similar to the mechanism of STDS stabilization by non-dissipative confinement along the longitudinal ($t$) coordinate [53], which corresponds to an additive phase mode-locking in a laser, the dissipative confinement due to synchronous pulse pumping or active amplitude mode-locking leads to a stable STDS, as well. Such stabilization is possible at very low values of the modulation “depth” parameter $\omega=G_{0}/\mathscr{T}^{2}\approx\Lambda/\mathscr{T}^{2}\approx 10^{-4}$. This means that gain variations along a laser axis could be very small. For instance, it means that a pumping pulse, synchronized with the cavity round-trip, could be very “long” compared to the STDS temporal width. The contribution of dynamical gain saturation lowers this value to $\approx 10^{-6}$. Simultaneously, there is a lower threshold on the value of the saturated net gain $|\Lambda^{\prime}|$. In our case, it was $\approx 0.01$. Too low net gain could not provide the 3D dissipative confinement and, thereby, STDS stabilization. We also found that the most optimal stabilization region is located near $\kappa\approx|\Lambda^{\prime}|/2$. Considering the dynamical gain saturation, the corresponding value is $\kappa\approx\Lambda/4-\Lambda/3$. It is important that spectral dissipation ($\tau$) is required for STDS stabilization. One may assume that the too-low value of this parameter cannot prevent a tendency to the STDS collapse.
The numerical simulations demonstrated that, besides the STDS destabilization due to decay for low $|\Lambda^{\prime}|$, $\kappa\rightarrow 0\land|\Lambda^{\prime}|$, and small $\omega$ and $\tau$, there is a tendency to STDS splitting in the time domain with the growth of $\kappa$ if $\omega$ is not sufficiently high. That means a “transverse” confinement needs an adequate “longitudinal” one. Despite the case of longitudinal phase confinement [53], we did not observe complicated spatial dynamics and splitting that testifies to the intensification of the mode cleaning process. Although “ring-like” spatial structures and the STDS “breathing” in the transverse direction are possible in the initial stages of evolution.
@@
We found that the gain saturation could significantly impact the STDS dynamics. For the considered values of gain coefficient $G_{ss}$ there is a confined range of driving frequency $\omega$ providing the STDS stability, and this region is the broadest in the vicinity of $\kappa\approx 0.0003-0.0004$. The main mechanism of destabilization is STDS decay. A possible explanation is the following. Too small $\omega$ does not provide sufficient confinement along $t-$dimension for a chosen $G_{ss}$. Too large driven frequency squeezes a gain window so STDS cannot be amplified sufficiently to start a Kerr effect for spatial squeezing, and spatial confinement becomes destructive. Also, the dynamical gain saturation defines some maximum $\kappa$ because it decreases an effective gain, and the spatial losses lead to the STDS decay. STDS peak power and width oscillations appear for small $\kappa$. These oscillations can be extremely strong and accompanied by alternating between asymmetric one and two pulses.
@@
We must note, that the cavity round-trip period could excess essentially $2L_{w}/c$ in order to provide the STDS energy scalability [59]. This method is practiced in high-energy passively mode-locked oscillators (e.g., see [12, 30]). In a waveguide oscillator, it is possible to use the Herriott cell or fiber loop for this aim. The last approach could be the most effective for the STDS generation in a fiber laser [53].
@@
In conclusion, we propose a mechanism of STDS stabilization using longitudinally and transversely graded gain in a waveguide laser with transversely graded refractive index and Kerr nonlinearity. Practically, that means realizing a DKLM in a waveguide laser operating in ADR and pumped by a Gaussian beam with slow amplitude temporal modulation synchronized with a laser repetition rate. In some sense, that is a dissipative analog of “pancake-like” confining potential in a weakly dissipative BEC. The analytical solutions for such STDS are obtained in the framework of VA depending on the parameters of the pump beam, such as its width, modulation depth, and saturated net gain. The essential ingredient of our model is spectral dissipation due to finite spectral gain bandwidth, which contributes to STDS stabilization. Numerical simulations under cylindrical symmetry conditions demonstrate the STDS’s stability within a limited region of parameters defined by net gain, pump beam width, and spectral bandwidth. Considering the dynamical gain saturation, we defined the main destabilization scenarios: decay and STDS oscillations asymmetrically alternating one and two pulses. We assume that the self-control of the transverse modes by the combination of enhanced nonlinearity, specific for the distributed nonlinear systems like waveguides composed from highly-nonlinear media, with the dissipative confinement, could provide the breakthrough in developing the new generation of high-stable and energy scalable sources of STDS. In principle, this approach is a further development of the DKLM concept that proved its efficiency for energy-scalable solid-state lasers. Moreover, it could be useful for stabilizing other localized coherent structures, e.g., a weakly dissipative BEC.
Funding
@@
The Norwegian Research Council projects #303347 (UNLOCK), #326503 (MIR), #32641 (Lammo-3D), and ALTA Lasers AS supported this work.

### v2

@@ 3 Discussion and Conclusion
Similar to the mechanism of STDS stabilization by non-dissipative confinement along the longitudinal ($t$) coordinate Kalashnikov and Wabnitz (2022), which corresponds to an active phase mode-locking in a laser, the dissipative confinement due to synchronous pulse pumping or active amplitude mode-locking leads to a stable STDS, as well. Such stabilization is possible at very low values of the modulation “depth” parameter $\omega=G_{0}/\mathscr{T}^{2}\approx\Lambda/\mathscr{T}^{2}~{}10^{-6}$. For instance, it means that a pumping pulse, synchronized with the cavity round-trip, could be “long” compared to the STDS temporal width.For instance, the normalizations presented in the Tables 2,3 result in the dimensional $\mathscr{T}\approx 8$ ps for $\Lambda=0.1$ , 10 cm ZnS-waveguide, $|\beta_{2}|=$2000 fs2/cm, and $\omega=10^{-6}$.
The linear stability analysis based on the variational approximation demonstrates that the stable STDS exists within confined regions of $\kappa$, $\Lambda^{\prime}$, and $\omega$ parameters. The numerical simulations demonstrate that the main destabilization scenario is an exponential STDS decay. Also, there is a tendency to STDS splitting in the time domain with the growth of $\kappa$ and $\omega$. For such a scenario, the “rogue wave” dynamics is characteristic. At variance with the case of longitudinal phase confinement Kalashnikov and Wabnitz (2022), we did not observe complicated spatial dynamics and splitting that testifies to the intensification of the mode cleaning process. It is important that spectral dissipation ($\tau$) is required for STDS stabilization. One may assume that the too-low value of this parameter cannot prevent a tendency to the STDS collapse.
@@
We found that the gain saturation could significantly impact the STDS dynamics. For the considered values of gain coefficient $G_{ss}$ there is a confined range of driving frequency $\omega$ providing the STDS stability, and this region is the broadest in the vicinity of $\kappa\approx\Lambda/4-\Lambda/3$. The primary mechanism of destabilization is STDS decay. A possible explanation is the following. Too small $\omega$ does not provide sufficient confinement along $t-$dimension for a chosen $G_{ss}$. Too large driven frequency squeezes a gain window so STDS cannot be amplified sufficiently to start a Kerr effect for spatial squeezing, and spatial confinement becomes destructive. Also, the dynamic gain saturation defines some maximum $\kappa$ because it decreases an effective gain, and the spatial losses lead to the STDS decay. The STDS peak power and width oscillations appear with the $\kappa$-decrease. These oscillations can be extremely strong and accompanied by alternating between asymmetric one and two pulses.
@@
We must note that the cavity round-trip period could excess essentially $2L_{w}/c$ to provide the STDS energy scalability Demesh et al. (2023). This method is practiced in high-energy passively mode-locked oscillators (e.g., see Südmeyer et al. (2008); Zhang et al. (2022)). In a waveguide oscillator, it is possible to use the Herriott cell or fiber loop for this aim. The last approach could be the most effective for the STDS generation in a fiber laser Kalashnikov and Wabnitz (2022).
@@
In conclusion, we propose a mechanism of STDS stabilization using longitudinally and transversely graded gain in a waveguide laser with transversely graded refractive index and Kerr nonlinearity. Practically, that means realizing a DKLM in a waveguide laser operating in ADR and pumped by a Gaussian beam with slow amplitude temporal modulation synchronized with a laser repetition rate. In some sense, that is a dissipative analog of “pancake-like” confining potential in a weakly dissipative BEC. The analytical solutions for such STDS are obtained in the framework of VA depending on the parameters of the pump beam, such as its width, modulation depth, and saturated net gain. The essential ingredient of our model is spectral dissipation due to finite spectral gain bandwidth, which contributes to STDS stabilization. Numerical simulations under cylindrical symmetry conditions demonstrate the STDS’s stability within a limited region of parameters defined by net gain, pump beam width, and spectral bandwidth. Considering the dynamic gain saturation, we defined the main destabilization scenarios: decay and STDS oscillations asymmetrically alternating one and two pulses. We assume that the self-control of the transverse modes by the combination of enhanced nonlinearity, specific for the distributed nonlinear systems like waveguides composed from highly-nonlinear media, with the dissipative confinement, could provide the breakthrough in developing the new generation of high-stable and energy scalable sources of STDS. In principle, this approach is a further development of the DKLM concept that proved its efficiency for energy-scalable solid-state lasers. Moreover, it could be useful for stabilizing other localized coherent structures, e.g., a weakly dissipative BEC.
Acknowledgments
@@
The authors would like to thank the Norwegian Research Council projects #303347 (UNLOCK), #326503 (MIR), #32641 (Lammo-3D), and ALTA Lasers AS supported this work. We are grateful to anonymous reviewers for the valuable comments that allowed us to make the essential improvements in the presentation of results.
Appendix A Euler-Lagrange-Kantorovich equations
The substitution of the ansatz (7) into the reduced Lagrangian (6) with a subsequent integration results in the following Euler-Lagrange equations (left-hand side of Eq(5)) for the ansatz parameters (a prime denotes $d/dZ$):
The substitution of (7) into the right-hand side of Eq. (5) leads to the Euler-Lagrange-Kantorovich equations:
One may obtain the equations for the evolution of the ansatz parameters from (12):
The steady-state points of (13) give Eqs. (8,9).
Appendix B Soliton spectrum
To find the Fourier image of (7), we use the fact that the chirp $\chi\ll 1$ in an anomalous dispersion regime. This fact results from calculations of (8,9). Then, it is possible to use a Fourier image expansion over $\chi$:
$\mathcal{F}\{\psi(r,t)\}(\Omega,\varpi)\approx\alpha\sqrt{\frac{2\pi\rho^{2}}{1-2i\theta\rho^{2}}}e^{-\frac{\Omega^{2}\rho^{2}}{2(1-2i\theta\rho^{2})}}\times$
[…]
The integral evaluation leads to:
Appendix C Jacobian
The calculated Jacobian for the system (13) is
where $\theta_{0}$, $\chi_{0}$, $\rho_{0}$, $\alpha_{0}$, and $T_{0}$ are the stationary points of the system (14).